OpenStudy (anonymous):

solve using elimination 9x + 2y = 2 21x + 6y = 4

OpenStudy (anonymous):

-3 ( 9x + 2y = 2 ) 21x + 6y = 4 results in -18x -6y = -6 21x +6y = 4 Which you then combined the terms and simplify -18x + 21x = 3x -6y + 6y = 0y -6 + 4 = -2 so we have: 3x +0y = -2 or x=-2/3 Plugging a known x into either equation will lead us to what y equals: 21(-2/3) + 6y = 4 -14 +6y = 4 6y = 4 +14 6y = 18 y=18/6 = 3 y=3 and x= -2/3

OpenStudy (anonymous):

why did you times it by -3 at the beginning?

OpenStudy (texaschic101):

the reason that you multiplied by a -3 was to eliminate the y's in the problem.

OpenStudy (anonymous):

so you just pick a number to illuminate one of them?

OpenStudy (anonymous):

Yes, it doesnt matter which equation you multiply nor does it matter which variable you eliminate first.

OpenStudy (anonymous):

how do you know what number will illuminate?

OpenStudy (raden_zaikaria):